OpenStudy (anonymous):

What is the equation of the line that contains the point (–5, –1) and has a slope of 4? Write in slope-intercept form. A. y = 4x – 19 B. y = 4x + 19 C. y = 4x – 1 D. y = 4x + 1

OpenStudy (anonymous):

In the standard form we have : y=mx+d So in there we have m = 4 ( m means slope ) and use this formula to find this :contains the point (–5, –1): y-y2=m(x-x2) x2=-5 y2=-1 and x and y are in the standard form . Got it ?:)

OpenStudy (anonymous):